# Transformers for dynamical systems learn transfer operators in-context

### v1

@@ Abstract
Large-scale foundation models for scientific machine learning adapt to physical settings unseen during training, such as zero-shot transfer between turbulent scales. This phenomenon, in-context learning, challenges conventional understanding of learning and adaptation in physical systems. Here, we study in-context learning of dynamical systems in a minimal setting: we train a small two-layer, single-head transformer to forecast one dynamical system, and then evaluate its ability to forecast a different dynamical system without retraining. We discover an early tradeoff in training between in-distribution and out-of-distribution performance, which manifests as a secondary double descent phenomenon. We discover that attention models apply a transfer-operator forecasting strategy in-context. They (1) lift low-dimensional time series using delay embedding, to detect the system’s higher-dimensional dynamical manifold, and (2) identify and forecast long-lived invariant sets that characterize the global flow on this manifold. Our results clarify the mechanism enabling large pretrained models to forecast unseen physical systems at test without retraining, and they illustrate the unique ability of attention-based models to leverage global attractor information in service of short-term forecasts.
@@
Scientific foundation models are pretrained on vast amounts of data, with a goal of discovering generalizable patterns applicable to unseen problem domains. Animating such efforts is the observation that, at sufficient scale, these models exhibit the emergent capability to solve unseen tasks during testing [33]. For example, models trained on low-resolution turbulent snapshots exhibit the spontaneous ability to forecast higher-resolution grids [24, 29, 19]. In some cases, models trained on certain classes of PDEs, like diffusion equations, develop the ability to forecast new classes of PDEs, like wave equations [23]. These findings align with similar observations for large language models, which exhibit the emergent ability to solve unseen tasks, like novel logical puzzles or mathematical derivations, at sufficient scale [7].
@@
The underlying mechanism, in-context learning, differs from classical machine learning approaches that dichotomize learning during training from testing during inference. Instead, pretraining provides a sufficiently large model with vast background experience and an underlying world model, allowing it to quickly adapt to new tasks based on minimal context [37, 2]. In traditional scientific machine learning, a trained model acts as a surrogate for the dynamical process that generated the training data, and so a trained model can, at best, only generate test outputs matching the distribution of the training data, known as in-distribution (ID) generalization. In contrast, in-context learning enables out-of-distribution (OOD) generalization on new data distributions [17]. However, the underlying mechanisms enabling in-context learning and OOD generalization in SciML are poorly understood, even as recent empirical studies have uncovered robust signatures of generalization in areas as diverse as turbulence closure modelling, large eddy convection, or magnetohydrodynamics [11, 32].
@@
A probabilistic forecast model $\hat{p}(x_{t+1}|\mathbf{x}_{C})\in\mathbb{R}^{V}$ accepts the context vector $\mathbf{x}_{C}\in\mathbb{Z}^{C}$ comprising $C$ past tokens, and returns a vector $\hat{p}(x_{t+1})\in\mathbb{R}^{V}$ corresponding to next-token probabilities across the vocabulary. Following prior works on minimal transformers, we model $\hat{p}$ using a minimal GPT-like transformer, comprising two layers with a single attention head each, and relative positional encoding [7] (Appendix B). For each input context $\mathbf{x}_{C}$, transformers compute self-attention, comprising $C^{2}$ pairwise comparisons among all tokens in the context, resulting in attention weights $A(\mathbf{x}_{C})\in\mathbb{R}^{C\times C}$ that quantify the relative importance of each past token in influencing the next-token probabilities.
@@
Remarkably, we observe that, in this minimal setting, small pretrained transformers exhibit the ability to predict unseen dynamical systems (Fig. 1A). These forecasts exceed the quality of naive baselines like mean regression, suggesting that transformers employ a more sophisticated in-context forecast strategy by leveraging their uniquely long context. During training on a single system (Train-ID), we see an early underfitting-to-overfitting transition on Test-ID, consistent with the classical bias-variance tradeoff: the model initially learns generalizable information about the system’s dynamics, leading to improvements on both seen and unseen data (Fig. 1B, gray and blue curves). However, as the Train-ID loss continues to fall, the model begins to overfit, leading to an increase in Test-ID error. At even later stages of training, we observe epoch-wise double descent, a recently-characterized phenomenon in which the Test-ID error further falls again below the classical bias-variance optimum [5, 31, 25]. However, we discover a second, earlier double descent curve in the Test-OOD error (Fig. 1B, magenta curve). This finding implies that information generalizable to other dynamical systems becomes available to the transformer during the pretraining process, but that it is bounded from above by ID-system-specific information.
@@
Figure 1: Double descent during in-context learning of dynamical systems. (A) Forecasts from a transformer trained on univariate time series from one dynamical system (Train-ID) then evaluated on its ability to forecast unseen trajectories from the same system (Test-ID, blue), versus forecasts of trajectories from an unseen system (Test-OOD, magenta). Gray curve shows the context, a subset of the total training data. (B) Loss curves on the training data, versus held-out in-distribution and out-of-distribution validation sets. Ranges are standard errors across $100$ replicate experiments, each corresponding to a different randomly-sampled train/test pair of dynamical systems.
@@
Where does this generalization signal come from? We sample the transformer’s time-delayed conditional distribution $\hat{p}(x_{t+1}|x_{t-k})$, $k\leq C$ averaged across Test-OOD for all lag values $k\in 0,1,...,C$. We find that this distribution qualitatively resembles the true attractor of Test-OOD, implying that the univariate model identifies relevant information about the unobserved multivariate attractor during testing (Fig. 2A). This result is consistent with Takens’ theorem: a univariate time series arising from an attractor with intrinsic dimension $d_{\mathcal{M}}$ can be embedded using $d_{E}=2d_{\mathcal{M}}+1$ time delays. We thus refer to the time delay coordinates $\hat{\mathbf{y}}_{i}=[\mathbf{x}_{i-d_{E}},...,\mathbf{x}_{i}]$ as an estimate of the unobserved true phase space coordinates $\mathbf{y}$ [34, 6].
@@
To test the scaling predicted by Takens’ theorem, we generate new Test-OOD trajectories using the Lorenz-96 system, a dynamical system with a controllable number of dynamical variables $D$ [21]. This system’s manifold dimension $d_{\mathcal{M}}$ scales approximately linearly with $D$ (Appendix D). As we vary the number of dynamical variables in Test-OOD D=3,4,25D=3,4,...25, we evaluate the average Test-OOD accuracy of transformers originally trained on a different dynamical system (Train-ID) with $D=3$. For each test case, we compare the transformer’s empirical order-$k$ conditional distribution (averaged across Test-OOD) to a $k$-order Markov chain $p(x_{t+1}|x_{t},x_{t-1},...,x_{t-k-1})$ trained directly on Test-OOD, in order to estimate the effective order $k$ of the Markov chain best-approximating the transformer’s Test-OOD dynamics. We find that the order of the best-approximating Markov chain scales with the intrinsic dimension of Test-OOD (Fig. 2C). This finding is striking, because the univariate Test-OOD series contains no direct information about the true dimensionality $D$ or $d_{\mathcal{M}}$, suggesting that the transformer instead leverages information from time delay embedding in-context.
To confirm that this effect arises from the transformer’s internal representations, for each Test-OOD example, we calculate the attention rollout matrix $A_{\text{roll}}\in\mathbb{R}^{C\times C}$ [1]. This metric measures the sparsity of the model’s attention to different token positions, averaged across the entire testing set and both attention heads. We compute the stable rank of the rollout matrix ($d_{\text{latent}}=\text{srank}(A_{\text{roll}})$) as a measure of the intrinsic dimensionality of the transformer’s latent representations. We find that increasing attractor dimension $d_{\mathcal{M}}$ strongly correlates with increasing latent dimensionality $d_{\text{latent}}$. Because the univariate time series presented to the transformer lack direct information about dimensionality, we conclude that the trained model infers information about the underlying attractor via in-context learning. Thus, not only does the transformer embed the test time series, it does so adaptively by identifying when a univariate time series arises from a high-dimensional system.
@@
Figure 2: Transformers perform time-delay embedding during inference. (A) Empirical time-delayed next-token probabilities $p_{\text{model}}(x_{t+1}|x_{t-k})$ averaged across Test-OOD context for a transformer trained on a different system (Train-ID). (B) Exact next-token probabilities $p_{\text{true}}(x_{t+1}|x_{t-k})$ obtained from fitting a Markov chain on a long sample of Test-OOD. (C) The order of the closest-approximating Markov chain, versus the true intrinsic dimension $d_{\mathcal{M}}$ of the full attractor from which the univariate Test-OOD trajectory originates. (D) The effective dimension (stable rank) of the transformer’s attention rollout matrix versus the true intrinsic dimension of Test-OOD. Each point is an average over $100$ replicate models trained on a different low-dimensional dynamical system.
We next consider how the model pushes forward the dynamics on the dynamical coordinates $\hat{\mathbf{y}}$. Broadly, there are two ways that a surrogate model could represent a dynamical propagator, given finite-length context: (1) a differential form based on approximating time derivatives (as in multi-step integrators, or differential equation inference methods like SINDy), or (2) an integral approach that approximates a transfer operator (as in neural operators or discrete-time system identification) [9, 8]. Based on the coincidence of the trained model’s empirical logit distribution with a Markov chain, we hypothesize the latter: transformers approximate propagators in-context.
@@
Figure 3: Transformers learn in-context transfer operators on reconstructed dynamical manifolds. (A) (Top) The true invariant distribution ($\pi^{(0)}(\mathbf{y})$) and longest-lived metastable distributions ($\pi^{(i)}(\mathbf{y})$ $i>0$, almost-invariant sets) estimated from the fully-observed state space as the leading eigenvectors of the fully-observed transfer operator $p(\mathbf{y}_{t+1}|\mathbf{y}_{t})$ (ranked by $\left|\lambda_{i}\right|$). (Bottom) The invariant and metastable distributions of the transformer, based on its next-token prediction probabilities over time-delayed univariate Test-OOD input sequences. Distributions are estimated as the leading eigenvectors of the lag-$K$ conditional probabilities $\hat{p}(x_{t+1}...x_{t-K+2}|x_{t}...x_{t-K+1})$ derived from the transformer probabilities. Because the transformer is univariate, eigenvectors are plotted on the time-delay embedding $\hat{\mathbf{y}}\equiv[x_{t},...,x_{t-K}]$. (B) The eigenvalue spectrum of the transfer operator estimated from the fully-observed attractor, and the implicit transfer operator of the transformer. (C) The KL divergence between the ground-truth invariant distribution of the transfer operator, and transformer empirical operator’s distribution, versus both Test-OOD validation loss and training epoch (color). A 95% confidence interval linear fit is underlaid (blue).
@@
To test this hypothesis, we extract the empirical transition matrix of the transformer on the space of length-$k$ delay vectors $\hat{\mathbf{y}}_{t}$ by aggregating multi-step predictions from the model $\hat{p}(x_{t+1}\mid\mathbf{x}_{C})$. Briefly, across Test-OOD we tabulate each distinct $k$-gram ($\mathbf{x}_{k}$, $k<C$), and we group together all context sequences $\mathbf{x}_{C}$ that end in that $k$-gram. For each unique $k$-gram, we randomly-sample from all matching context sequences and predict the frequency of the next token $x_{t+1}$. By repeating this process separately for each $k$-gram, we can estimate a transfer operator on the time-delayed sequence $p(\mathbf{\hat{y}}_{t+1}\mid\mathbf{\hat{y}}_{t})\equiv p(x_{t+1},x_{t},...,x_{t-k}|x_{t},x_{t-1},...,x_{t-k-1})$. This row-stochastic transition matrix describes the model-implied dynamics on the lag-$k$ state space embedding of the univariate Test-OOD time series.
@@
We compare the empirical operator to the ground truth operator $p(\mathbf{y}_{t+1}\mid\mathbf{y}_{t})$, which we estimate from the fully-observed state space using Ulam’s method, which discretizes the full multivariate phase space and tabulates all transition rates between pairs of regions [35]. We observe that the leading eigenvalues (ranked by $\left|\lambda_{i}\right|$) closely match, implying that the transformer captures the dominant timescales associated with the Perron-Frobenius operator propagating the dynamics (Fig. 3B). The corresponding eigenfunctions represent the stationary distribution $\pi^{(0)}$ and subleading eigenfunctions $\pi^{(1)}$, $\pi^{(2)}$ corresponding to long-lived metastable trapping regions, known as almost-invariant sets [14]. These eigenfunctions occur in pairs due to symmetry (Fig. 3A).
To confirm that the onset of OOD generalization during training arises from test-time inference of the underlying propagator, we directly compare the Test-OOD loss to the KL divergence $D_{KL}(\pi^{(0)}\parallel\hat{\pi}^{(0)})$ between the true and model-implied invariant distributions. Across hundreds of unique pairs of systems, we observe consistent correlation between Test-OOD loss and distributional distance (Fig. 3C). We thus conclude that, as the transformer’s out-of-distribution performance improves, it learns in-context an increasingly accurate estimate of the Perron-Frobenius operator of the fully-observed true system.
@@
Taken together, our results show that, during inference, transformers trained on dynamical systems first learn to perform time delay embedding on the context, and then use co-occurrence statistics to approximate the true transfer operator of the fully-observed state space. Our work thus provides an example of how a learning model can extract a system-independent forecasting strategy despite system-specific training, beyond naive strategies observed in prior works such as parroting sequences from the context [39]. Our work thus provides a potential explanation for recent works showing the surprising ability of pretrained foundation models to forecast dynamical systems outside their training data, such as turbulent flows, weather fronts, or plasma dynamics [24, 27, 32]. Additionally, our observation of in-context time delay embedding supports findings throughout the broader time series literature, that univariate models often match the performance of multivariate models, even for datasets with strong covariates [26]. It also accords with a recent theoretical result showing that single-layer transformers can implement adaptive time-delay embedding in-context, given sufficient latent dimensionality [12]. Finally, our results support recent neural scaling laws showing improved performance of time series and scientific foundation models with additional context [20]. Reminiscent of similar findings in language models [38], our work demonstrates how the unique quadratic complexity of attention-based models enables the emergence of nontrivial strategies for estimating and propagating dynamical states.

### v2

@@ Abstract
Large-scale foundation models for scientific machine learning adapt to physical settings unseen during training, such as zero-shot transfer between turbulent scales. This phenomenon, in-context learning, challenges conventional understanding of learning and adaptation in physical systems. Here, we study in-context learning of dynamical systems in a minimal setting: we train a small two-layer, single-head transformer to forecast one dynamical system, and then evaluate its ability to forecast a different dynamical system without retraining. We discover an early tradeoff in training between in-distribution and out-of-distribution performance, which manifests as a secondary double descent phenomenon. We discover that attention models apply a transfer-operator forecasting strategy in-context. They (1) lift low-dimensional time series using delay embedding, to detect the system’s higher-dimensional dynamical manifold, and (2) identify and forecast long-lived invariant sets that characterize the global flow on this manifold. Our results clarify the mechanism enabling large pretrained models to forecast unseen physical systems at test time without retraining, and they illustrate the unique ability of attention-based models to leverage global attractor information in service of short-term forecasts.
@@
Scientific foundation models are pretrained on vast amounts of data, with a goal of discovering generalizable patterns applicable to unseen problem domains. Animating such efforts is the observation that, at sufficient scale, these models exhibit the emergent capability to solve unseen tasks during testing [32]. For example, models trained on low-resolution snapshots of turbulence exhibit the spontaneous ability to forecast flows on higher-resolution grids [23, 27, 18]. In some cases, models trained on certain classes of PDEs, like diffusion equations, develop the ability to forecast new PDEs, like wave equations [22]. These findings align with similar observations for large language models, which exhibit the emergent ability to solve unseen tasks, like logical puzzles or mathematical derivations, at sufficient scale [7].
@@
The underlying mechanism, in-context learning, differs from classical machine learning approaches that dichotomize learning during training from inference during testing. Instead, pretraining provides a sufficiently large model with vast background experience and, in some cases, a mechanistic world model, allowing it to adapt to new tasks based on minimal context [35, 2]. In traditional scientific machine learning, a trained model acts as a surrogate for the dynamical process that generated the training data, and so a trained model can, at best, only generate test outputs matching the distribution of the training data, known as in-distribution (ID) generalization. In contrast, in-context learning enables out-of-distribution (OOD) generalization on new data distributions [17]. However, the underlying mechanisms enabling in-context learning and OOD generalization in SciML are poorly understood, even as recent empirical studies have uncovered robust signatures of generalization in areas as diverse as turbulence closure modelling, molecular dynamics, and magnetohydrodynamics [10, 31, 30].
@@
A probabilistic forecast model $\hat{p}(x_{t+1}|\mathbf{x}_{C})\in\mathbb{R}^{V}$ accepts the context vector $\mathbf{x}_{C}\equiv x_{t},x_{t-1},\ldots,x_{t-C+1}\in\mathbb{Z}^{C}$ comprising $C$ past tokens, and returns a vector $\hat{p}\in\mathbb{R}^{V}$ corresponding to next-token probabilities across the vocabulary. We estimate $\hat{p}$ from data using a minimal GPT-like transformer, comprising two layers with a single attention head each, and relative positional encoding [7, 15, 28, 13] (Appendix B). For each input context $\mathbf{x}_{C}$, transformers compute self-attention, comprising $C^{2}$ pairwise comparisons among all tokens in the context, resulting in attention weights $A(\mathbf{x}_{C})\in\mathbb{R}^{C\times C}$ that quantify the relative importance of each past token in influencing the next-token probabilities.
@@
Remarkably, we observe that, in this minimal setting, small pretrained transformers exhibit the ability to predict unseen dynamical systems (Fig. 1A). These forecasts outperform naive baselines like mean regression, suggesting that transformers employ a more sophisticated in-context forecast strategy by leveraging their uniquely long context. During training on a single system (Train-ID), we see an early underfitting-to-overfitting transition on Test-ID, consistent with the classical bias-variance tradeoff: the model initially learns generalizable information about the system’s dynamics, leading to improvements on both seen and unseen data (Fig. 1B, gray and blue curves). However, as the Train-ID loss continues to fall, the model begins to overfit, leading to an increase in Test-ID error. At even later stages of training, we observe epoch-wise double descent, a recently-characterized phenomenon in which the Test-ID error further falls again below the classical bias-variance optimum [4, 29, 24]. However, we discover a second, earlier double descent curve in the Test-OOD error (Fig. 1B, magenta curve). This finding implies that information generalizable to other dynamical systems becomes available to the transformer during the pretraining process, but that it is bounded from above by ID-system-specific information.
@@
Figure 1: Double descent during in-context learning of dynamical systems. (A) Forecasts from a transformer trained on univariate time series from one dynamical system (Train-ID) then evaluated on its ability to forecast unseen trajectories from the same system (Test-ID, blue), versus forecasts of trajectories from an unseen system (Test-OOD, magenta). Gray curve shows the context, a subset of the total test data. (B) Loss curves on the training data, versus held-out in-distribution and out-of-distribution validation sets. Ranges are standard errors across $100$ replicate experiments, each corresponding to a different randomly-sampled train/test pair of dynamical systems.
[…]
$\mathcal{L}_{\text{OOD}}\approx\mathcal{H}\!\left(p^{\text{OOD}}_{k\text{-markov}}\right)+D_{\mathrm{KL}}\!\left(p^{\text{OOD}}_{k\text{-markov}}\,\middle\|\,p^{\text{ID}}_{k\text{-markov}}\right)$ (1)
This equation thus relates the error of the transformer on a new dynamical system to the degree of train/test shift.
When sufficient data is available the first term in Eq. 1 plateaus in $k$. However, finite context and training data limit the accuracy of the Markov description at large $k$ and so, for our experimental setting, we find that the bigram ($k=2$) case exhibits the strongest relationship. Across our experiments, we find that the OOD loss is well-described by the expected linear scaling $\mathcal{L}_{\mathrm{OOD}}\propto D_{\mathrm{KL}}^{(2)}$ in Eq. 1 (Fig. 2A, Spearman’s $\rho=0.72\pm 0.02$, $500$ system pairs). Our observations align with domain-adaptation theory, which predicts that target risk depends on the task-relevant difference between the source and target distributions [5, 21].
Figure 2: Scaling laws for out-of-distribution generalization in dynamical systems. (A) The test error of the unseen system (cross-entropy, Test-OOD) versus the difference between the training and testing sets (KL divergence between the attractor of Test-OOD and Train-ID). (B) The error in the steady-state invariant distribution of the transformer dynamics relative to the true invariant distribution of Test-OOD, versus the context length $C$. Dashed line shows the expected scaling for an estimator that counts observed state transitions in-context [13]. Points indicate $500$ transformers trained and tested on distinct ID/OOD dynamical system pairs, and solid lines indicate linear fits with 99% confidence intervals.
@@
What strategies allow the transformer to generalize? We sample the transformer’s time-delayed conditional distribution $\hat{p}(x_{t+1}|x_{t-k})$, $k\leq C$ averaged across Test-OOD for all lag values $k\in 0,1,\ldots,C$. We find that this distribution qualitatively resembles the true attractor of Test-OOD, implying that the univariate model identifies relevant information about the unobserved multivariate attractor during testing (Fig. 3A). This result is consistent with Takens’ theorem: if a univariate time series arises from an attractor with intrinsic dimension $d_{\mathcal{M}}$, embedding the time series with at least $d_{E}=2d_{\mathcal{M}}+1$ time delays recovers the key geometric features of the true attractor [33]. We denote the time delay coordinates $\hat{\mathbf{y}}_{i}=[x_{i-d_{E}+1},...,x_{i}]^{\top}$ as estimates of the unobserved true phase space coordinates $\mathbf{y}$ [6].
@@
To test the scaling predicted by Takens’ theorem, we generate new Test-OOD trajectories using the Lorenz-96 system, a dynamical system with a controllable number of dynamical variables $D$ [20]. This system’s manifold dimension $d_{\mathcal{M}}$ scales approximately linearly with $D$ (Appendix E). As we vary the number of dynamical variables in Test-OOD $D=3,4,\ldots,25$, we evaluate the average Test-OOD accuracy of transformers originally trained on a different dynamical system (Train-ID) with $D=3$. For each test case, we compare the transformer’s empirical $k$-order conditional distribution (averaged across Test-OOD) to a $k$-order Markov chain $p_{\text{k-markov}}^{\text{OOD}}(x_{t+1}|\mathbf{x}_{k})$ trained directly on Test-OOD, in order to estimate the effective order $k$ of the Markov chain best-approximating the transformer’s Test-OOD dynamics. We find that the order of the best-approximating Markov chain scales linearly with the intrinsic dimension of Test-OOD (Fig. 3C), consistent with Takens’ theorem. This finding is striking, because the univariate Test-OOD series contains no direct information about the true dimensionality $D$ or $d_{\mathcal{M}}$, suggesting that the transformer instead adapts during inference by estimating the intrinsic dimensionality of the input time series.
@@
Figure 3: Transformers perform time-delay embedding during inference. (A) Empirical time-delayed next-token probabilities $\hat{p}(x_{t+1}|x_{t-k})$ averaged across Test-OOD context for a transformer trained on a different system (Train-ID). (B) Exact next-token probabilities $p_{\text{true}}(x_{t+1}|x_{t-k})$ obtained from fitting a Markov chain on a long sample of Test-OOD. (C) The order of the closest-approximating Markov chain, versus the true intrinsic dimension $d_{\mathcal{M}}$ of the full attractor from which the univariate Test-OOD trajectory originates. (D) The effective dimension (stable rank) of the transformer’s attention rollout matrix versus the true intrinsic dimension of Test-OOD. Stable rank is scaled by the maximum possible rank. Annotated slopes indicate the linear scaling predicted by Takens’ theorem. Each point is an average over $100$ replicate models trained on a different low-dimensional dynamical system.
To confirm that this effect arises from the transformer’s internal representations, we calculate the attention rollout matrix $A_{\text{roll}}\in\mathbb{R}^{C\times C}$ [1]. This metric measures the context-dependent response of the model to different token positions, averaged across the entire testing set and both attention heads. We compute the stable rank of the rollout matrix ($d_{\text{latent}}=\text{srank}(A_{\text{roll}})$) as a measure of the intrinsic dimensionality of the transformer’s latent representations. We find that increasing attractor dimension $d_{\mathcal{M}}$ linearly correlates with increasing latent dimensionality $d_{\text{latent}}$. Because the univariate time series presented to the transformer lack direct information about dimensionality, we conclude that the trained model infers information about the underlying attractor via in-context learning. Thus, not only does the transformer embed the test time series, it does so adaptively by identifying when a univariate time series arises from a high-dimensional system.
@@
Figure 4: Transformers learn in-context transfer operators on reconstructed dynamical manifolds. (A) The eigenvalue spectrum of the transfer operator estimated from the fully-observed Test-OOD attractor $p(\mathbf{y}_{t+1}|\mathbf{y}_{t})$, and the time-lagged transfer operator estimated by sampling contiguous length-$k$ sequences from the transformer $\hat{p}(\hat{\mathbf{y}}_{t+1}|\hat{\mathbf{y}}_{t})=\hat{p}(x_{t+1},x_{t},...x_{t-k}|x_{t},x_{t-1},...x_{t-k-1})$. (B) (Top) The invariant distribution $\pi^{(0)}(\mathbf{y})$ and longest-lived metastable distributions ($\pi^{(i)}(\mathbf{y})$, $i>0$) of the true transfer operator, ranked by $\left|\lambda_{i}\right|$. (Bottom) The invariant and metastable distributions of the transformer, based on its implied transfer operator $\hat{p}(\hat{\mathbf{y}}_{t}|\hat{\mathbf{y}}_{t-1})$. Because the transformer is univariate, eigenvectors are plotted on the time-delay embedding $\hat{\mathbf{y}}\equiv[x_{t},x_{t-1}]^{\top}$.
We next consider how the model propagates the dynamics after estimating the dynamical coordinates $\hat{\mathbf{y}}$. Broadly, there are two ways that a surrogate model could represent a dynamical propagator, given finite-length context: (1) a differential form based on approximating time derivatives (as in multi-step integrators, or differential equation inference methods like SINDy), or (2) an integral approach that approximates a transfer operator (as in neural operators or discrete-time system identification) [9, 8]. Based on the coincidence of the trained model’s empirical next-token distribution with a Markov chain, we hypothesize the latter: transformers approximate propagators in-context.
@@
To test this possibility, we extract the empirical transition matrix of the transformer on the space of length-$k$ delay vectors $\hat{\mathbf{y}}_{t}$ by aggregating multi-step predictions from the model $\hat{p}(x_{t+1}\mid\mathbf{x}_{C})$. Briefly, across Test-OOD we tabulate each distinct $k$-gram ($\mathbf{x}_{k}$, $k<C$), and we group together all context sequences $\mathbf{x}_{C}$ that end in that $k$-gram. For each unique $k$-gram, we randomly-sample from all matching context sequences and predict the frequency of the next token $x_{t+1}$. By repeating this process separately for each $k$-gram, we estimate a transfer operator on the time-delayed sequence $\hat{p}(\mathbf{\hat{y}}_{t+1}\mid\mathbf{\hat{y}}_{t})\equiv\hat{p}(x_{t+1},x_{t},...,x_{t-k}|x_{t},x_{t-1},...,x_{t-k-1})$. This row-stochastic transition matrix describes the model-implied dynamics on the lag-$k$ state space embedding of the univariate Test-OOD time series. In dynamical systems theory, this transition matrix represents a coarse-grained estimate of the Perron-Frobenius operator, which describes how dynamical systems evolve distributions of trajectories over time [11].
@@
We compare the empirical operator to the ground truth operator $p(\mathbf{y}_{t+1}\mid\mathbf{y}_{t})$, which we estimate from the fully-observed state space using Ulam’s method: we discretize the full multivariate phase space and tabulate all transition rates between pairs of regions [34]. We observe that the leading eigenvalues (ranked by $\left|\lambda_{i}\right|$) closely match, implying that the transformer captures the dominant timescales associated with the Perron-Frobenius operator propagating the dynamics (Fig. 4A). The corresponding eigenfunctions represent the stationary distribution $\pi^{(0)}$ and subleading eigenfunctions $\pi^{(1)}$, $\pi^{(2)}$ corresponding to long-lived metastable trapping regions, known as almost-invariant sets [14]. These eigenfunctions occur in pairs due to symmetry (Fig. 4B).
Our findings agree with recent studies showing that minimal, two-layer transformers as we consider here can implement $k$-gram counting circuits in-context [13], potentially enabling estimation of the transfer operator via Ulam’s method. Given an estimate $\hat{p}(\mathbf{y}_{t+1}|\mathbf{y}_{t})$ of the true transfer operator ${p}(\mathbf{y}_{t+1}|\mathbf{y}_{t})$, the KL divergence scales quadratically to leading order $D_{\text{KL}}(p||\hat{p})\sim\mathcal{O}((p-\hat{p})^{2})$ when the discrepancy is small. Because the number of expected counts of a given state $i$ increases linearly in the context $N_{i}\approx\pi_{i}^{(0)}C$, bounds from eigenvalue perturbation theory predict $D_{\text{KL}}(p||\hat{p})\sim\mathcal{O}(C^{-1})$ (Appendix G). We find that this scaling holds in practice across many models trained on distinct Train-ID/Test-OOD system pairs (Fig. 2B). We thus conclude that, as the transformer’s out-of-distribution performance improves, it learns in-context an increasingly accurate estimate of the Perron-Frobenius operator of the fully-observed true system.
@@
Taken together, our results show that, during inference, transformers trained on dynamical systems first learn to perform time delay embedding on the context, and then use co-occurrence statistics to approximate the true transfer operator of the fully-observed state space. Our work thus provides an example of how a learning model can extract a system-independent forecasting strategy despite system-specific training, beyond naive strategies observed in prior works such as parroting sequences from the context [37]. Our work thus provides a potential explanation for recent works showing the surprising ability of pretrained foundation models to forecast dynamical systems outside their training data, such as turbulent flows, weather fronts, or plasma dynamics [23, 26, 30]. Additionally, our observation of in-context time delay embedding supports findings throughout the broader time series literature, that univariate models often match the performance of multivariate models, even for datasets with strong covariation [25]. It also accords with a recent theoretical result showing that single-layer transformers can implement adaptive time-delay embedding in-context, given sufficient latent dimensionality [12]. Finally, our results support recent neural scaling laws showing improved performance of time series and scientific foundation models with additional context [19]. Reminiscent of similar findings in language models [36], our work demonstrates how the unique structure of attention-based models enables nontrivial strategies for estimating and propagating dynamical states.